OpenStudy (anonymous):

what values of x would make following expression undefined 45\(9-x^2)

jhonyy9 (jhonyy9):

do you know a fraction when is undefined ?

jhonyy9 (jhonyy9):

so when the denominator is equal zero ,ok ? so in this case when will be the denominator zero ,for what value of x ?

jhonyy9 (jhonyy9):

so you can calcule it than make equale zero the denominator 9-x^2 = 0 add to the both sides x^2 9 = x^2 make squarroot from both sides and will get x_1,2 = +/- 3 hope this will help you and this is understandably